archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Incident Category Backend Core Status CLOSED Percent Complete Details We detected a fiber cut between the NRA of HEL59 and Roubaix A ticket was opened investigations are in progress on their side There is no service interruption security was resumed

    Original URL path: http://status.ovh.es/?do=details&id=5206 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    CLOSED Percent Complete Details The NRA is down optical problem we are investigating Date Friday 09 August 2013 04 47AM Reason for closing Done Comments 1 Comment by OVH Friday 09 August 2013 00 52AM There are works with our

    Original URL path: http://status.ovh.es/?do=details&id=5194 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details Tonight at around 0 00 we will be updating the version of the NRA CAN59 optical devices The purpose is to the withstand Lambda 10G This operation is transparent so no downtime is expected Date Friday 09 August 2013 00 41AM Reason for closing Done Comments 2 Comment by OVH Friday 09 August 2013 00 41AM upgrading cards service Comment by OVH

    Original URL path: http://status.ovh.es/?do=details&id=5192 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Tonight around 0 00 we will update the version of the optical equipment for Bordeaux region The aim is to support 10G lambda This operation is transparent so we don t expect a cut off Date Tuesday 06 August 2013

    Original URL path: http://status.ovh.es/?do=details&id=5163 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    CLOSED Percent Complete Details There is an issue with the NRA RON34 on the SFR collect A ticket has been opened and investigations are in progress on their side Date Tuesday 06 August 2013 04 37AM Reason for closing Done Comments 1 Comment by OVH Thursday 01 August 2013 16 41PM There is a air conditioning fault on the Montpellier Rondelet FT site The rise in temperature on site has

    Original URL path: http://status.ovh.es/?do=details&id=5140 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    send a technician to gsw tomorrow morning currently our customers will keep going through th2 12 Comment by OVH Tuesday 30 July 2013 06 34AM gsw 12 n5 is up now we will keep monitoring Comment by OVH Tuesday 30 July 2013 08 25AM Now we have a problem on pair 10 gsw 10 and th2 10 we are reloading th2 10 Comment by OVH Tuesday 30 July 2013 08 25AM Everything is OK following the reboot Comment by OVH Tuesday 30 July 2013 18 35PM Due to the optical problem we lost securisation between th2 12 n5 and gsw 12 n5 The customers continue to function on th2 12 n5 Comment by OVH Wednesday 31 July 2013 01 57AM We are now facing repetition issues on several optics DWDM equipping our DSL switches in Paris and used to connect our Globalswitch POPs and TH2 This has a several times impact on the traffic many cutoffs from few seconds to several minutes depending on customers We have therefore decided to change the techno and move the links GSW TH2 in DWDM active ADVA on a first half the concerned links This has to allow the stability of the infrastructure and

    Original URL path: http://status.ovh.es/?do=details&id=5126 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    departements currently recorded are 41 45 51 77 78 95 A ticket has been opened investigations are in progress on their side Date Wednesday 31 July 2013 19 12PM Reason for closing Done Comments 3 Comment by OVH Wednesday 31 July 2013 14 59PM The issue is with our collect operator s router in Valenton Comment by OVH Wednesday 31 July 2013 17 45PM Progressive connection increase from 3 50

    Original URL path: http://status.ovh.es/?do=details&id=5133 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details We have an electrical issue in the NRA we are looking for the origin of the problem Date Wednesday 31 July 2013 09 38AM Reason for closing Done Comments 3 Comment by OVH Monday 29 July 2013 00 23AM The energy room was flooded following the storm tonight Comment by OVH Monday 29 July 2013 00 23AM DSLAM is up now Comment

    Original URL path: http://status.ovh.es/?do=details&id=5110 (2015-10-15)
    Open archived version from archive



  •